# Litmus tests of the flat ${\Lambda\text{CDM}}$ model and model-independent measurement of $H_{0}r_{\text{d}}$ with LSST and DESI

### v1

@@ Abstract
In this analysis we apply a model-independent framework to test the flat $\Lambda$CDM cosmology using simulated SNIa data from the upcoming Large Synoptic Survey Telescope (LSST) and combined with simulated Dark Energy Spectroscopic Instrument (DESI) five-years Baryon Acoustic Oscillations (BAO) data. We adopt an iterative smoothing technique to reconstruct the expansion history from SNIa data, which, when combined with BAO measurements, facilitates a comprehensive test of the Universe’s curvature and the nature of dark energy. The analysis is conducted under three different mock true cosmologies: a flat $\Lambda$CDM universe, a universe with a notable curvature ($\Omega_{k,0}=0.1$), and one with dynamically evolving dark energy. Each cosmology demonstrates different kinds and varying degrees of deviation from the standard model predictions. We forecast that our reconstruction technique can constrain cosmological parameters, such as the curvature ($\Omega_{k,0}$) and $H_{0}r_{\text{d}}$, with a precision of approximately 0.52% for $H_{0}r_{\text{d}}$ and between 0.032 to 0.037 for $\Omega_{k,0}$, competitive with current cosmic microwave background constraints, without assuming any form of dark energy.
@@ 1 Introduction
The concordance cosmological ${\Lambda\text{CDM}}$ model relies on a few assumptions. Among them, homogeneity and isotropy lead to the Friedman-Robertson-Lemaître-Walker (FLRW) metric. The discovery of cosmic acceleration led to the establishment of the concordance ${\Lambda\text{CDM}}$ model. In this model, the energy budget of the Universe is dominated by a cosmological constant $\Lambda$, and matter by cold dark matter (CDM). The most stringent constraints on the model have been obtained by cosmic microwave background (CMB) measurements from the Planck satellite [1], baryon acoustic oscillations (BAO) [2, 3, 4, 5, 6] and Type Ia Supernovae (SNIa) [7, 8, 9, 10, 11]. CMB constrains the curvature parameter to be $\Omega_{k}=-0.044^{+0.018}_{-0.015}$. Most of these constraints however are direct fits to the data assuming a particular model, typically the ${\Lambda\text{CDM}}$ model or an extension. While this leverages the power of Bayesian statistics to provide tight constraints on the model parameters, it is important to also test the validity of the underlying hypotheses.
As opposed to model-fitting, model-independent approaches provide a healthy check on the validity of a given model. Due to their extra flexibility compared to model-fitting approaches, they may also reveal the presence of unexpected features in the data and may be used to look for unaccounted for systematics [12, 13, 14].
The rise of tensions, in particular the Hubble tension, and the unknown nature of the main components of the model are strong hints for cosmologist to search for evidence for beyond-${\Lambda\text{CDM}}$ physics [15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Recent data from the Dark Energy Spectroscopic Instrument (DESI) showed evidence for a phantom dark energy or a dynamical dark energy [5]. This was further confirmed by studying physics-driven modifications of ${\Lambda\text{CDM}}$ [24] as well as data-driven reconstructions of the expansion history [25].
@@
The curvature of the Universe is another fundamental question. The Planck 2018 data are consistent with a flat universe but do not exclude a small negative curvature [1], see also [26]. DESI DR1 finds a preference for a small positive curvature, which disappears when combined with Planck.
@@
In this work, we used a model-independent approach, namely, the iterative smoothing method, to reconstruct the expansion history from SNIa and combine it with distance measurements from BAO to perform a litmus test of the pillars of the ${\Lambda\text{CDM}}$ model, the FLRW metric, the flatness of the Universe, and the cosmological constant, as well to constrain $H_{0}r_{\text{d}}/c$.
@@ 2.1 Fiducial Models
To show the potential of our model-independent approach, we simulate data for three different cosmologies:
  • a flat-${\Lambda\text{CDM}}$ universe, with $\Omega_{\text{m},0}=0.315$ and a constant equation of state $w=-1$.
  • a curved ${\Lambda\text{CDM}}$ universe, with ($\Omega_{\text{m},0},\Omega_{k,0})=(0.315,0.1)$. While such a large curvature is consistent with Pantheon+ or DESI DR1 data, it is excluded by Planck 2018 data. However, it is nonetheless interesting to see the effect of curvature and to apply our curvature test to a curved universe.
  • a flat universe with dynamical dark energy. We chose the phenomenologically emergent dark energy (PEDE) model [28], which has no extra degree of freedom relative to ${\Lambda\text{CDM}}$. The dark energy density evolves as
@@
We note that the data and covariance matrix were initially generated for the flat ${\Lambda\text{CDM}}$ case, and we reused the same noise and covariance matrix for the other two cosmologies. While for a proper forecast, one should apply the pipeline from [29] using the other two fiducial cosmologies, we assume here that the resulting noise and covariance should be affected minimally by the change of fiducial cosmology. On the other hand, the mock BAO data are all generated with Gaussian white noise.
Table 1 summarizes our three fiducial cosmologies.
Fiducial Cosmology $\Omega_{\text{m},0}$ $h$ $\Omega_{k,0}$ $w$
flat-${\Lambda\text{CDM}}$ 0.315 0.685 0 -1
$\Omega_{k,0}=0.1$ 0.315 0.685 0.1 -1
PEDE 0.315 0.685 0 Eq. (2.1b)
Table 1: Fiducial cosmologies used in this work
@@ 2.2 LSST SNIa Simulations
2.2.1 Vera C. Rubin Observatory (The LSST)
The forthcoming LSST survey, scheduled to commence in 2025, represents the most comprehensive optical survey planned to date, marking a significant milestone for its generation. This survey will run for a decade and will extensively explore the southern hemisphere’s sky through a large, ground-based, wide-field observatory utilising six optical passband filters. This combination is designed to achieve an unparalleled balance of depth, coverage area, and observational frequency. Over the course of the survey, the LSST is expected to catalogue millions of supernovae [30]. Operating from the Rubin Observatory, the Simonyi Survey Telescope features an $8.4$ m mirror (with an effective aperture of $6.7$ m) and is equipped with a state-of-the-art $3200$-megapixel camera, providing a $9.6$ deg2 field of view. Approximately $90\%$ of the telescope’s time will be dedicated to a deep-wide-fast survey mode, systematically covering an area of $18,000$ square degrees roughly $800$ times across all bands over a decade, resulting in a co-added map with a depth of $r-27.5$ (called as the Wide Fast Deep (WFD) observing strategy in the LSST). This extensive data collection is projected to amass around $32$ trillion observations of $20$ billion galaxies and an equivalent number of stars, primarily supporting the survey’s main scientific objectives [30]. The remaining $10\%$ of the time will be reserved for specific initiatives, such as the Very Deep and Very Fast time-domain surveys  [^2] Deep Drilling Fields (DDF), which are still in the planning stages.
@@ 2.2.2 Mock Data Generation
This study employs simulated SNIa data to examine dark energy results from the forthcoming LSST observations, as detailed in [29]. The supernova data are generated through the LSST Dark Energy Science Collaboration (DESC) Time Domain (TD) pipeline and the SNANA software [31]. The procedure encompasses four primary stages, depicted in Figure 4 of [29].
@@ 2.3 Mock BAO data from DESI 5-year
The Dark Energy Spectroscopic Instrument (DESI) is a stage-IV experiment aiming to constrain dark energy by mapping the Universe in three dimensions [6].
BAO are another probe of the expansion history from large-scale structures. Strictly speaking, the transverse and radial modes of the BAO constrain
$\frac{d_{\text{A}}(z)}{r_{\text{d}}}$ $=\frac{c}{H_{0}r_{\text{d}}}\frac{\mathcal{D}(z)}{1+z},\text{and}$ (2.3a)
$\frac{c}{H(z)r_{\text{d}}}$ $=\frac{c}{H_{0}r_{\text{d}}}\frac{1}{h(z)},$ (2.3b)
where
and $r_{\text{d}}$ is the sound horizon at the drag epoch $z_{\mathrm{d}}$.
@@ 3 Method
describe the time evolution of the dark energy density, $\Omega_{\text{m},0}$ and $\Omega_{k,0}$ are the matter and curvature density parameters today respectively, and $w=p/\rho$ is the equation of state of dark energy. For the cosmological constant $\Lambda$, $w=-1$ and $f_{\text{DE}}(z)\equiv 1$.
The dimensionless comoving distance is defined by
$\mathcal{D}(z)$ $=\int_{0}^{z}\frac{\mathrm{d}x}{h(x)}.$ (3.3)
$(1+z)d_{\text{A}}(z)$ $=\frac{d_{\text{L}}(z)}{1+z}=d_{\text{C}}(z)=\frac{c}{H_{0}}\mathcal{D}(z).$ (3.4)
The distance modulus from supernovae is
[…]
Several litmus tests have been put forward to test various aspects of the ${\Lambda\text{CDM}}$ model. In particular, in an FLRW universe, we have [37, 38, 39]
$\mathcal{O}_{k}(z)$ $=\frac{\Theta^{2}(z)-1}{\mathcal{D}^{2}(z)}=\Omega_{k,0},$ (3.6a)
where
$\Theta$ $=h\mathcal{D}^{\prime}.$ (3.6b)
In a flat-FLRW universe, $\Theta(z)\equiv 1$.
The $O_{m}$ diagnostic, introduced by [40], is defined as
$O_{m}(z)$ $=\frac{h^{2}(z)-1}{(1+z)^{3}-1}=\Omega_{\text{m},0}$ (3.7)
in a flat-${\Lambda\text{CDM}}$ universe. These two diagnostics are litmus tests respectively of the FLRW metric (and its curvature) and of the flat-${\Lambda\text{CDM}}$ model. Departure from $O_{m}=$constant, or from $\mathcal{O}_{k}=$ constant are signs of departure from flat-${\Lambda\text{CDM}}$ and FLRW respectively.
@@
$q$ $=-\frac{1}{H}\frac{\ddot{a}}{a}=-(1+z)\frac{h^{\prime}}{h}-1.$ (3.8)
$q>0$ indicate a decelerating Universe, while $q<0$ for an accelerating universe.
@@ 3.1 Curvature test
The transverse and radial BAO modes respectively give us ${d_{\text{A}}}/{r_{\text{d}}}=d_{\text{C}}/((1+z)r_{\text{d}})$ and $c/{(Hr_{\text{d}})}$, while the supernovae provide $\mathcal{D}$ and $\mathcal{D}^{\prime}$, which are independent of both $H_{0}$ and $r_{\text{d}}$. We can therefore combine these two measurements to estimate $H_{0}r_{\text{d}}$ in two independent ways:
$\frac{H_{0}r_{\text{d}}}{c}$ $=\mathcal{D}\frac{r_{\text{d}}}{d_{\text{C}}}\mbox{, and}$ (3.9a)
$\frac{H_{0}r_{\text{d}}}{c}$ $=\frac{1}{h}\frac{Hr_{\text{d}}}{c}=\mathcal{D}^{\prime}\frac{Hr_{\text{d}}}{c},$ (3.9b)
where the last equality holds only for a flat-FLRW universe. Therefore, we can combine the BAO observables with $\mathcal{D}$ and $\mathcal{D}^{\prime}$ measured by SNIa to constrain $H_{0}r_{\text{d}}$. One way to do this is to reconstruct a smooth comoving distance $\mathcal{D}(z)$, its derivative $\mathcal{D}^{\prime}(z)$, and expansion history $h(z)$ from SNIa at any arbitrary redshift $z$, and evaluate those curves at the BAO redshifts $z_{\text{BAO}}$. To do the reconstructions, we apply the iterative smoothing method to simulated LSST SNIa as described in Section 3.2. We then simply divide the central value of the BAO measurements $d_{\text{C}}/r_{\rm d}$ and $d_{\text{H}}/r_{\rm d}$ by the value of the reconstructed $\mathcal{D}(z_{\text{BAO}})$, $\mathcal{D}^{\prime}(z_{\text{BAO}})$, and $h(z_{\rm BAO)}$, and then propagate the BAO error. The results of this calculation are seen in Fig. 4.
@@
$\Theta$ $=h\mathcal{D}^{\prime}=\frac{Hr_{\text{d}}}{H_{0}r_{\text{d}}}\mathcal{D}^{\prime}=\frac{Hr_{\text{d}}}{c}\frac{d_{\text{C}}}{r_{\text{d}}}\frac{\mathcal{D}^{\prime}}{\mathcal{D}}$ (3.10)
and $\mathcal{O}_{k}$ only with the BAO and the smooth reconstructions from the SNIa data. Similarly to $H_{0}r_{\rm d}$, we evaluate $\Theta(z)$ by multiplying the central values of the BAO data ($\frac{Hr_{\text{d}}}{c}$ and $\frac{d_{\text{C}}}{r_{\text{d}}}$) with the smooth functions reconstructed from the SNIa ($\frac{\mathcal{D}^{\prime}(z_{\rm BAO})}{\mathcal{D}(z_{\rm BAO})}$), evaluated at the BAO redshifts.
Figure 1: Results of the SNIa smoothing colour-coded by $\Delta\chi^{2}=\chi^{2}-\chi^{2}_{\Lambda\text{CDM}}$ for the flat-${\Lambda\text{CDM}}$ fiducial cosmology. The top-left panel shows the smooth residuals. The middle-left panel shows $\mathcal{D}/\mathcal{D}_{\text{fid}}$, the bottom-left panel shows $h_{\text{fid}}/h$, the middle-right panel shows $O_{m}$, and the bottom-right panel $q$. The middle- and bottom-left panels also show the DESI-like BAO mock data normalized by the fiducial cosmology.
Figure 2: Same as Fig. 1, fiducial model: ${\Lambda\text{CDM}}$ $(\Omega_{k,0}=0.1)$ Note that in the bottom-left panel, the BAO data are $h_{\text{fid}}/h$, but the smooth SNIa reconstructions show $\mathcal{D}^{\prime}/\mathcal{D}^{\prime}_{\text{fid}}$.
Figure 3: Same as Fig. 1, fiducial model: PEDE

### v3

@@ Abstract
In this analysis we apply a model-independent framework to test the flat $\Lambda$CDM cosmology using simulated SNIa data from the upcoming Legacy Survey of Space and Time (LSST) and combined with simulated Dark Energy Spectroscopic Instrument (DESI) five-years Baryon Acoustic Oscillations (BAO) data. We adopt an iterative smoothing technique to reconstruct the expansion history from SNIa data, which, when combined with BAO measurements, facilitates a comprehensive test of the Universe’s curvature and the nature of dark energy. The analysis is conducted under four different mock true cosmologies: Two curvatures ($\Omega_{k,0}=0$ and 0.1) and two models of dark energy: a cosmological constant $\Lambda$ and the phenomenologically emergent dark energy. We forecast that our reconstruction technique can constrain cosmological parameters, such as the curvature ($\Omega_{k,0}$) and $c/H_{0}r_{\text{d}}$, with spread due to the SNIa uncertainties up to ±$\pm 4\%$ and ±$\pm 0.1$ respectively, without assuming any form of dark energy.
@@ 1 Introduction
The concordance cosmological ${\Lambda\text{CDM}}$ model relies on a few assumptions. Among them, homogeneity and isotropy lead to the Friedman-Robertson-Lemaître-Walker (FLRW) metric. The discovery of cosmic acceleration led to the establishment of the concordance ${\Lambda\text{CDM}}$ model. In this model, the energy budget of the Universe today is dominated by a cosmological constant $\Lambda$, and matter by cold dark matter (CDM). The most stringent constraints on the model have been obtained by cosmic microwave background (CMB) measurements from the Planck satellite [1], baryon acoustic oscillations (BAO) [2, 3, 4, 5, 6] and Type Ia Supernovae (SNIa) [7, 8, 9, 10, 11].
@@
The rise of tensions, in particular the Hubble tension, and the unknown nature of the main components of the model are strong hints for cosmologists to search for evidence for beyond-${\Lambda\text{CDM}}$ physics [12, 13, 14, 15, 16, 17, 18, 19, 20]. While the simplest explanation for the late-time cosmic acceleration is dark energy in the form of a cosmological constant, a dynamical DE is not excluded [21]. Recent BAO data from the Dark Energy Spectroscopic Instrument (DESI) showed evidence for a phantom dark energy or a dynamical dark energy [5]. This was further confirmed by studying physics-driven modifications of ${\Lambda\text{CDM}}$ [22] as well as data-driven reconstructions of the expansion history [23].
@@
Another fundamental question is the curvature of the Universe, summarized in the current curvature density parameter $\Omega_{k,0}$. The Planck 2018 (P18) data favor a negative curvature $-0.095<\Omega_{k,0}<-0.007$ at 99% [1, 24, 25], However, combining P18 with external data yields results consistent with a flat universe [26, 27]. Interestingly, DESI DR1 finds a preference for a small positive curvature, which disappears when combined with Planck.
Most of these constraints however are direct fits to the data assuming a particular model, typically the ${\Lambda\text{CDM}}$ model or an extension. While this leverages the power of Bayesian statistics to provide tight constraints on the model parameters, it is important to also test the validity of the underlying hypotheses. As opposed to model-fitting, model-independent approaches provide a healthy check on the validity of a given model. Due to their extra flexibility compared to model-fitting approaches, they may also reveal the presence of unexpected features in the data and may be used to look for unaccounted for systematics [28, 29, 30].
@@
In this work, we used a model-independent approach, namely, the iterative smoothing method, to reconstruct the expansion history from SNIa and combine it with distance measurements from BAO to perform litmus tests of the pillars of the ${\Lambda\text{CDM}}$ model, the FLRW metric, the flatness of the Universe, and the cosmological constant, as well as to constrain $c/H_{0}r_{\text{d}}$. Specifically, our goal is twofolds: we aim to quantify the constraining power of the combination of LSST SNIa and DESI Y5 BAO to (1) test the FLRW metric, and (2) to measure the curvature parameter and $H_{0}r_{\text{d}}$ independently of the nature of DE.
2 Theory
where
describes the time evolution of the dark energy density, $\Omega_{\text{m},0}$ and $\Omega_{k,0}$ are the matter and curvature density parameters today respectively, and $w=p/\rho$ is the equation of state of dark energy. For the cosmological constant $\Lambda$, $w=-1$ and $f_{\text{DE}}(z)\equiv 1$.
The dimensionless comoving distance is defined by
$\mathcal{D}(z)$ $=\int_{0}^{z}\frac{\mathrm{d}x}{h(x)}.$ (2.3)
$(1+z)d_{\text{A}}(z)$ $=\frac{d_{\text{L}}(z)}{1+z}=d_{\text{M}}(z)=\frac{c}{H_{0}}\mathcal{D}(z).$ (2.4)
The distance modulus from supernovae is
[…]
BAO, on the other hand, measure both the distances and expansion. Strictly speaking, the transverse and radial modes of the BAO give
$\frac{d_{\text{A}}(z)}{r_{\text{d}}}$ $=\frac{c}{H_{0}r_{\text{d}}}\frac{\mathcal{D}(z)}{1+z}\mbox{, and}$ (2.6a)
$\frac{c}{H(z)r_{\text{d}}}$ $=\frac{c}{H_{0}r_{\text{d}}}\frac{1}{h(z)}\mbox{, where}$ (2.6b)
and $r_{\text{d}}$ is the sound horizon at the drag epoch $z_{\mathrm{d}}$. $H_{0}r_{\text{d}}$ is thus an interesting parameter combining early- and late-time physics, which may play an important role in the $H_{0}$ tension [31, 32, 33, 34].
@@ 3.1 Fiducial Models
To show the potential of our model-independent approach, we simulate data for four different fiducial cosmologies. All fiducial cosmologies share the same $(\Omega_{\text{m},0},h)=(0.315,0.7)$, which is the cosmology used in [36].
  • Flat-$\Lambda$: a flat-${\Lambda\text{CDM}}$ universe, with a constant equation of state $w=-1$.
  • $k$-$\Lambda$: a curved ${\Lambda\text{CDM}}$ universe, with $\Omega_{k,0}=0.1$. While such a large curvature is consistent with Pantheon+ or DESI DR1 data, it is excluded by Planck 2018 data. However, it is nonetheless interesting to see the effect of curvature and to apply our curvature test to a curved universe.
  • Flat-PEDE: a flat universe with dynamical dark energy. We chose the phenomenologically emergent dark energy (PEDE) model [37], which has no extra degree of freedom relative to ${\Lambda\text{CDM}}$. The dark energy density evolves as
@@
  • $k$-PEDE: a curved universe with $\Omega_{k,0}=0.1$ and PEDE.
We note that the SNIa mock data and covariance matrix were initially generated for the flat ${\Lambda\text{CDM}}$ case. We extracted the noise $\mu-\mu_{\text{fid}}$ and applied it to the other three fiducial cosmologies. Although for a proper forecast, one should apply the pipeline from [36] to the other three fiducial cosmologies, we assume here that the resulting noise and covariance should be affected minimally by the change of fiducial cosmology. On the other hand, the mock BAO data are all generated with a Gaussian white noise as detailed in Section 3.3.
Table 1 summarizes our four fiducial cosmologies.
Fiducial Cosmology $\Omega_{k,0}$ $w$ $\chi^{2}_{{\Lambda\text{CDM}}}$ $\Delta\chi^{2}_{\mathrm{fid}}$
${\Lambda\text{CDM}}$ 0 -1 7.48 1.27
$k$-${\Lambda\text{CDM}}$ 0.1 -1 11.68 -2.93
PEDE 0 Eq. (3.1b) 8.21 0.54
$k$-PEDE 0.1 Eq. (3.1b) 7.55 1.20
Table 1: Fiducial cosmologies used in this work. The fourth column shows the ${\Lambda\text{CDM}}$ $\chi^{2}_{lcdm}$ to the data, used as a reference throughout this work, and the last column shows the $\Delta\chi^{2}$ of the fiducial model with respect to the best-fit ${\Lambda\text{CDM}}$.
@@ 3.2 LSST SNIa Simulations
3.2.1 The Rubin Observatory Legacy Survey of Space and Time (LSST)
The forthcoming LSST survey, scheduled to commence in 2025, represents the most comprehensive optical survey planned to date, marking a significant milestone for its generation. This survey will run for a decade and will extensively explore the southern hemisphere’s sky through a large, ground-based, wide-field observatory utilising six optical passband filters. This combination is designed to achieve an unparalleled balance of depth, coverage area, and observational frequency. Over the course of the survey, the LSST is expected to catalogue millions of supernovae [38]. Operating from the Vera C. Rubin Observatory, the Simonyi Survey Telescope features an $8.4$ m mirror (with an effective aperture of $6.7$ m) and is equipped with a state-of-the-art $3200$-megapixel camera, providing a $9.6$ deg2 field of view. Approximately $90\%$ of the telescope’s time will be dedicated to a deep-wide-fast survey mode, systematically covering an area of $18,000$ square degrees roughly $800$ times across all bands over a decade, resulting in a co-added map with a depth of $r-27.5$ (called as the Wide Fast Deep (WFD) observing strategy in the LSST). This extensive data collection is projected to amass around $32$ trillion observations of $20$ billion galaxies and an equivalent number of stars, primarily supporting the survey’s main scientific objectives [38]. The remaining $10\%$ of the time will be reserved for specific initiatives, such as the Very Deep and Very Fast time-domain surveys, [^2] Deep Drilling Fields (DDF) which are still in the planning stages.
@@ 3.2.2 Mock Data Generation
This study employs simulated SNIa data to examine dark energy results from the forthcoming LSST observations, as detailed in [36]. The supernova data are generated through the LSST Dark Energy Science Collaboration (DESC) Time Domain (TD) pipeline and the SNANA software [39]. In this analysis, our SN data refers to using the corresponding Hubble diagram, and the covariance matrix provided by [36]. The Hubble diagram consists of the redshift ($z$) of measurements, the corresponding luminosity distance ($D_{L}$), and their statistical plus systematic errors combined ($\sigma(D_{L})$). The procedure encompasses four primary stages, depicted in Figure 4 of [36].
@@
We would also like to mention that the SN analysis done using the LSST Time-Domain pipeline used a Gaussian SN likelihood, but the method presented in our analysis here, can as well incorporate asymmetric and non-Gaussian errors. Therefore our methodology will still apply, even if we had a different covariance matrix and likelihood for our data (including non-gaussian likelihoods). The fourth column of Table 1 shows the ${\Lambda\text{CDM}}$ $\chi^{2}$ to the SNIa data for each fiducial cosmology. As expected, the flat-$\Lambda$ cosmology has the lowest $\chi^{2}_{{\Lambda\text{CDM}}}$. The last column shows $\Delta\chi^{2}=\chi^{2}_{\text{fid}}-\chi^{2}_{{\Lambda\text{CDM}}}$, the difference between the $\chi^{2}$ of the fiducial cosmology to that of the ${\Lambda\text{CDM}}$ best-fit. It is worth noting that, since we use the same noise for all cosmologies, $\chi^{2}_{\text{fid}}$ is the same for all four cosmologies, in this case 8.75. Therefore, $\Delta\chi^{2}_{\text{fid}}$ is larger for the flat-$\Lambda$ cosmology.
@@ 3.3 Mock BAO data from DESI 5-year
The Dark Energy Spectroscopic Instrument (DESI) is a stage-IV experiment aiming to constrain dark energy by mapping the Universe in three dimensions [6]. In particular, DESI is measuring the BAO, a probe of the expansion history from large-scale structures.
We generated DESI 5-year mock BAO, assuming a Gaussian white noise where the errors on $d_{\text{A}}$ and $H$ are taken from [35], assuming a correlation coefficient of 0.4 between $d_{\text{A}}$ and $H$, and then propagated to errors on $d_{\text{A}}$ and $d_{\text{H}}$ via the usual error propagation formula for correlated variables.
@@ 4 Method
Several litmus tests have been put forward to test various aspects of the ${\Lambda\text{CDM}}$ model. In particular, the $\mathcal{O}_{k}$ diagnostic [45, 46, 31] is defined as
$\mathcal{O}_{k}(z)$ $=\frac{\Theta^{2}(z)-1}{\mathcal{D}^{2}(z)},$ (4.1a)
@@
$\Theta(z)$ $=h(z)\mathcal{D}^{\prime}(z),$ (4.1b)
where denotes the derivative with respect to redshift. In an FLRW universe, $\mathcal{O}_{k}(z)\equiv\Omega_{k,0}$, and $\Theta(z)=\sqrt{1+\Omega_{k,0}\mathcal{D}^{2}}$, and in a flat-FLRW universe, $\Theta(z)\equiv 1$. It is important to realize that these equalities only assume the FLRW metric, and in particular, are independent of the dark energy evolution.
The $\mathcal{O}m$ diagnostic, introduced by [47], is defined as
$\mathcal{O}m(z)$ $=\frac{h^{2}(z)-1}{(1+z)^{3}-1}=\Omega_{\text{m},0}$ (4.2)
in a flat-${\Lambda\text{CDM}}$ universe. These two diagnostics are litmus tests respectively of the FLRW metric (and its curvature) and of the flat-${\Lambda\text{CDM}}$ model. Departure from $\mathcal{O}m=$ constant, or from $\mathcal{O}_{k}=$ constant are signs of departure from flat-${\Lambda\text{CDM}}$ and FLRW respectively.
@@
$q$ $=-\frac{1}{H}\frac{\ddot{a}}{a}=-(1+z)\frac{h^{\prime}}{h}-1,$ (4.3)
where $\dot{}$ denotes the derivative with respect to time $t$. $q>0$ indicate a decelerating Universe, while $q<0$ for an accelerating universe.
@@ 4.1 Curvature test
The transverse and radial BAO modes respectively give us ${d_{\text{A}}}/{r_{\text{d}}}=d_{\text{M}}/((1+z)r_{\text{d}})$ and $d_{\text{H}}/r_{\text{d}}=c/{(Hr_{\text{d}})}$, while the supernovae provide $\mathcal{D}$ and $\mathcal{D}^{\prime}$, which are independent of both $H_{0}$ and $r_{\text{d}}$. Intuitively, the SNIa constrain (the log of) $\mathcal{D}$, and we can obtain $\mathcal{D}^{\prime}$ by calculating the smooth derivative of $\mathcal{D}$(see Section 4.2). We can therefore combine these two measurements to estimate $H_{0}r_{\text{d}}$ in two independent ways:
$\frac{c}{H_{0}r_{\text{d}}}$ $=\frac{1}{\mathcal{D}(z)}\frac{d_{\text{M}}(z)}{r_{\text{d}}}\mbox{, and}$ (4.4a)
$\frac{c}{H_{0}r_{\text{d}}}$ $=h(z)\frac{c}{Hr_{\text{d}}}=\frac{1}{\mathcal{D}^{\prime}(z)}\frac{d_{\text{H}}(z)}{r_{\text{d}}},$ (4.4b)
where the last equality in eq. (4.4b) holds only for a flat-FLRW universe. Therefore, we can combine the BAO observables with $\mathcal{D}$ and $\mathcal{D}^{\prime}$ measured by SNIa to constrain $H_{0}r_{\text{d}}$. One way to do this is to reconstruct a smooth comoving distance $\mathcal{D}(z)$, its derivative $\mathcal{D}^{\prime}(z)$, and expansion history $h(z)$ from SNIa at any arbitrary redshift $z$, and evaluate those curves at the BAO redshifts $z_{\mathrm{BAO}}$. To do the reconstructions, we apply the iterative smoothing method to simulated LSST SNIa as described in Section 4.2. We then simply divide the central value of the BAO measurements $d_{\text{M}}/r_{\rm d}$ and $d_{\text{H}}/r_{\rm d}$ by the value of the reconstructed ${\mathcal{D}}$ and $\mathcal{D}^{\prime}$ evaluated at the BAO redshifts and then propagate the BAO error.
@@
$\Theta(z)$ $=\frac{H(z)r_{\text{d}}}{H_{0}r_{\text{d}}}\mathcal{D}^{\prime}(z)=\frac{d_{\text{M}}(z)/r_{\text{d}}}{d_{\text{H}}(z)/r_{\text{d}}}\frac{\mathcal{D}^{\prime}(z)}{\mathcal{D}(z)}$ (4.5)
and $\mathcal{O}_{k}$ only with the BAO and the smooth reconstructions from the SNIa data. Similarly to $H_{0}r_{\rm d}$, we evaluate $\Theta(z)$ by multiplying the central values of the BAO data ($d_{\text{H}}/r_{\text{d}}$ and ${d_{\text{M}}}/{r_{\text{d}}}$) with the smooth functions reconstructed from the SNIa ($\mathcal{D}^{\prime}/\mathcal{D}$), evaluated at the BAO redshifts $z_{\rm BAO}$, and propagating the error by taking the covariance between $d_{\text{M}}$ and $d_{\text{H}}$ as detailed in Appendix B.
Figure 1: Results of the SNIa smoothing colour-coded by $\Delta\chi^{2}_{\text{SN},n}=\chi^{2}_{\text{SN},n}-\chi^{2}_{\text{SN, ${\Lambda\text{CDM}}$}}$ for the flat-${\Lambda\text{CDM}}$ fiducial cosmology. The top-left panel shows the smooth residuals. The middle-left panel shows $\mathcal{D}/\mathcal{D}_{\text{fid}}$, the bottom-left panel shows $h_{\text{fid}}/h$, the middle-right panel shows $\mathcal{O}m$, and the bottom-right panel $q$. The middle- and bottom-left panels also show the DESI-like BAO mock data normalized by the fiducial cosmology.
Figure 2: Same as Fig. 1, fiducial model: $k$-${\Lambda\text{CDM}}$ Note that in the bottom-left panel, the BAO data are $h_{\text{fid}}/h$, but the smooth SNIa reconstructions show $\mathcal{D}^{\prime}/\mathcal{D}^{\prime}_{\text{fid}}$.
Figure 3: Same as Fig. 1, fiducial model: Flat-PEDE
Figure 4: Same as Fig. 1, fiducial model: $k$-PEDE